with the current, a sailboat travels 9 km in 3 hrs. however, if it travels against the current, it takes 3 times to sail the same distance. how fast is the boat sailing on still water? how fast does the river flow?
Answer by checkley77(12844)

9/3=3 MPH DOWN STREAM.
9/(3*3)=9/9=1 KPH UPSTREAM.
(3-1)/2=2/2=1 KPH IS THE CURRENT.
PROOF:
3-1=2 KPH THE SPEED OF THE SAIL BOAT.
1+1=2 DITTO.